Graphically illustrate equilibrium Qd = 1800 - 10p and Qs = 200 + 10p
1
Expert's answer
2020-04-30T09:46:23-0400




"1800-10p=200+10p"


"20p=1600"


"p=80"


"Q=1800-10\times80=1000"